The volume of a box is 60um^3(micro)
If the length is 50 cm and the length is 3.5*(10^-3)km, what is the height, in pm?

Answer by Alan3354(69443)

The volume of a box is 60um^3(micro)
If the length is 50 cm and the length is 3.5*(10^-3)km, what is the height, in pm?
----------
Vol = L*W*H
60 um^3 = 50 cm * 3.5 meters * h = 50000 microns * 3500000 microns * h
h (microns) = 60/(5E4*3.5E6) = 60/1.75E11 microns
h = (60/1.75)E-5 pm (assuming pm = picometers)
h = 3.42857E-4 picometers
------------------
h =~ 342.857 attometers